40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 763     Entry time: Wed Jul 30 01:08:50 2008
Author: rana 
Type: Summary 
Category: SUS 
Subject: SUS Drift Screen 
This is a snap of the SUS Drift screen with all of the optics biases set back to their nominal
values except for the MC which Rob aligned and I didn't feel like mis-aligning. The reference
on the screen is from 3/25 when Andrey felt that Rob had a good IFO alignment.

Anything more than a few thousand is significant and more than 10k means something is wrong:



I wailed on the PRM for awhile and was able to loosen it up a little. The LL & LR sensors now
show some life one the dataviewer. The UL & UR are still railed ~1.6 V so that means that the
optic is pitched back. With aggressive pitch wailing I can see the PRM's ULR/UR sensors go
rail to rail so that means that the magnets are still on - although they may be half busted.
If they're OK we should be able to just re-sling this guy.


Did the same on SRM. The OSEM values have shifted on these, but not disastorously. The SIDE,
however, is completely unresponsive. The little signal I see when driving is is probably just
capacitive pickup in the cables. Have to vent to fix this one.


ITMX Has good life in all but the LR & UR channels. They respond, but the signal is very weak.
Seems like these magnets have not fallen off but that they are not between the LED/PD anymore.


ITMY seems ok. Check the spectra to be sure.


BS seems ok as well. Swings freely and no kinks in the swinging sensor waveforms. Check the spectra.
Attachment 1: infection-2.png  17 kB  | Hide | Hide all
infection-2.png
ELOG V3.1.3-